# Comment on ”Monochromatization interaction region optics design for direct s-channel Higgs production at FCC-ee”

### v1

@@ Abstract
The original article [1] can be logically divided into two parts: 1) the selection of main parameters for monochromatization and 2) interaction region optics design; the comment pertains only to the first part. The authors of [1] state that ”The purpose of this paper is to report on the development of realistic IR optics designs for monochromatization at the FCC-ee”. However, the proposed parameters do not seem very realistic and raise many questions.
@@ 1 Piwinski angle and collision overlap area
In general, the collision overlap area [3] (that is Ref. [14] in [1]) is expressed as $L_{i}=\sigma_{z}\big/\sqrt{1+\varphi^{2}}$. As the authors rightly noted, the $L_{i}\big/\beta_{y}^{*}$ ratio characterizes the hourglass effect. But only for $\varphi\gg 1$ the simplification of the form $L_{i}\approx 2\sigma_{x}^{*}\big/\theta_{c}$ is valid. However, the typical value of Piwinski angle in Tables 1 and 2 (with monochromatization) is $\varphi\approx 1$, so the exact expression for $L_{i}$ must be used. Accordingly, the line ”Overlap collision …” in the Tables should be corrected, which will also affect $R_{hg}$ and the calculated luminosity.
@@ 2 Synchrotron tune and $t\bar{t}$ optics
The modulation of the spin precession frequency by synchrotron oscillations [4] is characterized by the synchrotron modulation index: $\zeta=\nu_{0}\sigma_{\delta}\big/Q_{s}$, where $\nu_{0}\approx 142$ is the spin tune at 62.5 GeV and $\sigma_{\delta}\approx 5.5\cdot 10^{-4}$. For a reliable determination of the spin frequency, one should aim for $\zeta<1.5$, and in any case $\zeta<2$ must be satisfied. It follows that $Q_{s}$ should be at least greater than 0.04, and preferably greater than 0.05. In principle, this is possible in $t\bar{t}$ optics, but it will require raising the RF voltage from 0.17 to 1.01.21.0-1.2 GV. As a consequence, $\sigma_{z}$ will decrease several times, greatly increasing beamstrahlung and making this option less attractive. In any case, many of the values in Table 2 should change significantly.
@@ 3 Vertical dispersion
The time available for direct s-channel Higgs production experiment will be limited. Therefore, it is critical to have a sufficiently high luminosity to obtain the required number of events. The dependence of the Higgs event rate on luminosity and $\sigma_{W}$ can be expressed as $\dot{N}_{H}\propto L\Big/\sqrt{\Gamma_{H}^{2}+\sigma_{W}^{2}}$, where $\Gamma_{H}\approx 4.1$ MeV. The original idea was to increase $\dot{N}_{H}$ by monochromatization in the hope that the reduction in the luminosity would be small. Then the goals became more realistic: $L$ can decrease in approximately the same proportion as $\sigma_{W}$, so that $\dot{N}_{H}$ remains almost unchanged. But if monochromatization leads to a noticeable (several times) decrease in $\dot{N}_{H}$, then it loses its meaning.
With this in mind, let us consider the scheme with $D_{y}^{*}\neq 0$ again. On the one hand, it is not hindered by the horizontal crossing angle, and this is the main advantage. On the other hand, increasing $\sigma_{y}^{*}$ leads to a decrease in $L$ by a factor of $\lambda$ (see Eq. (2) in [1]), which is still within acceptable limits. But this works only without BS, that is, for a small bunch population and, consequently, at low luminosity. And taking BS into account, $\dot{N}_{H}$ inevitably falls several times. The advantage of horizontal dispersion over vertical one becomes clear if we take into account the fact that for flat beams, increasing $\sigma_{x}^{*}$ weakens BS, while increasing $\sigma_{y}^{*}$ does not. That is why the option with vertical dispersion was not even considered in the CDR [5].
The results presented in [1] also demonstrate a multiple reduction in $\dot{N}_{H}$ in the case of $D_{y}^{*}\neq 0$. Therefore, to avoid misleading readers, it would be useful to make a clear and distinct conclusion that such a scheme should be rejected.
@@ 4 Monochromatization factor
Eq. (3) in [1] is obviously incorrect. To understand this, it suffices to consider the limiting case $D_{y}^{*}=0$, $\varphi\gg 1$ and $D_{x}^{*}\sigma_{\delta}\big/\sqrt{\varepsilon_{x}\beta_{x}^{*}}\gg\varphi$. Then we get $\lambda\gg 1$, which cannot be the case for $\varphi\gg 1$ – this is clear from the geometry shown in Fig. 2 in [1]. Everything is explained simply: the authors took Eq. (3) from [2] (that is Ref. [13] in [1], see Eq. (17) there), but did not take into account that the Piwinski angle in [2] is defined incorrectly: instead of the full horizontal beam size $\sigma_{x}^{*}$, the betatron size $\sigma_{x\beta}^{*}=\sqrt{\varepsilon_{x}\beta_{x}^{*}}$ was substituted into the formula for determining $\varphi$; this issue is considered in detail in [3].
To correct Eq. (3), it is necessary to multiply $\varphi^{2}$ by the factor $(D_{x}^{*2}\sigma_{\delta}^{2}\big/(\varepsilon_{x}\beta_{x}^{*})+1)$. From this, in particular, it follows that $\sigma_{W}$ should increase from 26.4 to about 37.9 MeV for ZH4IP and from 20.2 to about 41.8 MeV for ZH2IP, making the latter scheme less attractive. These values do not agree well with the simulation results presented in Table 3, which raises questions about this simulation.
@@ 5 Vertical emittance
In Section 3.3, the authors reasonably noted that $\varepsilon_{y}$ was calculated with the assumed FCC-ee coupling ratio $\varepsilon_{y}\big/\varepsilon_{x}=0.2\%$. If $D_{x}^{*}\neq 0$, $\varepsilon_{x}$ will increase several times due to beamstrahlung, and it would be too optimistic to assume that $\varepsilon_{y}$ will remain unchanged. For example, $\varepsilon_{x}$ increased about 3.7 times for ZH4IP. Under the conservative assumption that about half of this increase is converted to $\varepsilon_{y}$ with the same $0.2\%$ ratio, the luminosity will decrease by approximately 1.5 times.
@@ 6 Parameter optimization
The Abstract of [1] contains a phrase that is perceived as a promise that will be fulfilled: ”This measurement is significantly facilitated if the CM energy spread of $e^{+}e^{-}$ collisions can be reduced to a level comparable to the natural width of the Higgs boson, $\Gamma_{H}$ = 4.1 MeV, without substantial loss in luminosity”. And here is the result: $\sigma_{W}$ became $\sim$38 MeV, while the luminosity decreases by 2.3 times (taking into account the aforementioned factor of 1.5). Is this what we were aiming for?
Obviously, to get $\lambda\gg 1$ in a collision scheme with crossing angle and without crab crossing, we need to make $\varphi\ll 1$. To achieve this, $\sigma_{z}$ can be reduced (for example, $\sim$1.5 times) by increasing the RF voltage and $D_{x}^{*}$ should be significantly increased. The statement about a ”minimum bunch spacing of 25 ns” also raises questions. Perhaps this limitation exists at low energy (Z-pole, 45.6 GeV), where the total beam current is 3.5 times higher. However, at 62.5 GeV things become easier, so decreasing $N_{b}$ (along with increasing the number of bunches $n_{b}$) can be a useful option.
An adequate example of parameter optimization can be found in the CDR [5], where the emphasis was on achieving the required $\sigma_{W}$ and luminosity values. Of course, much has changed since then, including the perimeter, the number of IPs, the arc cell type at low energies and hence $\alpha_{C}$, $\varepsilon_{x}$, etc. So, it will be necessary to re-optimize everything, and then, what will be the most difficult part, try to implement it in the IR optics design.

### v2

@@ Abstract
The original article [1] can be logically divided into two parts: 1) the selection of main parameters for monochromatization and 2) interaction region optics design; the comment pertains only to the first part. The authors of [1] state that ”The purpose of this paper is to report on the development of realistic IR optics designs for monochromatization at the FCC-ee”. However, the proposed parameters do not seem very realistic and raise many questions; due to space limitations, we will only consider the most important ones.
@@ 1 Piwinski angle and collision overlap area
In the general case, the length of interaction area [3] (that is Ref. [14] in [1]) is expressed as
$L_{i}=\frac{\sigma_{z}}{\sqrt{1+\varphi^{2}}}.$ (1)
For $\varphi\gg 1$ and $\theta_{c}\ll 1$, this can be simplified to:
$L_{i}\approx\frac{\sigma_{z}}{\varphi}\approx\frac{2\sigma_{x}}{\theta_{c}}.$ (2)
As the authors rightly noted, the $L_{i}\big/\beta_{y}^{*}$ ratio characterizes the hourglass effect. But using Eq. (2) for collision overlap area in [1] is unjustified, since the typical value of Piwinski angle in Tables 1 and 2 (with monochromatization) is $\varphi\approx 1$. The exact expression (1) for $L_{i}$ should be used instead. Accordingly, the line ”Overlap collision …” in the Tables should be corrected, which will also affect $R_{hg}$ and the calculated luminosity.
@@ 2 Synchrotron tune and $t\bar{t}$ optics
The modulation of the spin precession frequency by synchrotron oscillations [4] is characterized by the synchrotron modulation index: $\zeta=\nu_{0}\sigma_{\delta}\big/Q_{s}$, where $\nu_{0}\approx 142$ is the spin tune at 62.5 GeV and $\sigma_{\delta}\approx 5.5\cdot 10^{-4}$. For a reliable determination of the spin frequency, one should aim for $\zeta<1.5$, and in any case $\zeta<2$ must be satisfied. It follows that $Q_{s}$ should be at least greater than 0.04, and preferably greater than 0.05. In principle, this is possible in $t\bar{t}$ optics, but it will require raising the RF voltage from 0.17 to 1.01.21.0-1.2 GV. As a result, $\sigma_{z}$ will decrease several times, which will have both positive and negative consequences. Positive aspects include weakening the hourglass effect and reducing the Piwinski angle, which is necessary for monochromatization. Negative aspects include enhancement of collective instabilities for short bunches and increased beamstrahlung, which will lead to a significant increase in $\varepsilon_{x}$ and a decrease in $\lambda$. To avoid this, it will be necessary to reduce the bunch population several times (and, consequently, increase the number of bunches), which will lead to a decrease in luminosity. In any case, many of the values in Table 2 will have to change significantly, so in its current form, this Table does not make much sense.
@@ 3 Vertical dispersion
The time available for direct s-channel Higgs production experiment will be limited. Therefore, it is critical to have a sufficiently high luminosity to obtain the required number of events. The dependence of the Higgs event rate on luminosity and $\sigma_{W}$ can be expressed as
$\dot{N}_{H}\propto\frac{L}{\sqrt{\Gamma_{H}^{2}+\sigma_{W}^{2}}},$ (3)
where $\Gamma_{H}\approx 4.1$ MeV. The original idea was to increase $\dot{N}_{H}$ by monochromatization in the hope that the reduction in the luminosity would be small. Then the goals became more realistic: $L$ can decrease in approximately the same proportion as $\sigma_{W}$, so that $\dot{N}_{H}$ remains almost unchanged. But if monochromatization leads to a noticeable (several times) decrease in $\dot{N}_{H}$, then it loses its meaning.
With this in mind, let us consider the scheme with $D_{y}^{*}\neq 0$ again. On the one hand, it is not hindered by the horizontal crossing angle, and this is the main advantage. On the other hand, increasing $\sigma_{y}^{*}$ leads to a decrease in $L$ by a factor of $\lambda$, which is still within acceptable limits. But this works only without BS, that is, for a small bunch population and, consequently, at low luminosity. And taking BS into account, $\dot{N}_{H}$ will inevitably decrease many times over. That is why the option with vertical dispersion was not even considered in the CDR [5].
The advantage of horizontal dispersion over vertical one becomes clear if we take into account the fact that for flat beams, increasing $\sigma_{x}^{*}$ weakens BS, while increasing $\sigma_{y}^{*}$ does not. The results presented in [1] also demonstrate a multiple reduction in $\dot{N}_{H}$ in the case of $D_{y}^{*}\neq 0$. Therefore, to avoid misleading readers, it would be useful to make a clear and distinct conclusion that such a scheme should be rejected.
@@ 4 Monochromatization factor
Eq. (3) in [1] is obviously incorrect. To understand this, it suffices to consider the limiting case $D_{y}^{*}=0$, $\varphi\gg 1$ and $D_{x}^{*}\sigma_{\delta}\big/\sqrt{\varepsilon_{x}\beta_{x}^{*}}\gg\varphi$. Then we get $\lambda\gg 1$, which cannot be the case for $\varphi\gg 1$ – this is clear from the geometry shown in Fig. 2 in [1]. Everything is explained simply: the authors took Eq. (3) from [2] (that is Ref. [13] in [1]), see Eq. (17) there. However, they did not take into account that the Piwinski angle is defined incorrectly in [2]: instead of the full horizontal beam size $\sigma_{x}^{*}$, the betatron size $\sigma_{x\beta}^{*}=\sqrt{\varepsilon_{x}\beta_{x}^{*}}$ was substituted into the formula for determining $\varphi$; this issue is considered in detail in [3].
The correct expression for $\lambda$ in the case $D_{y}^{*}=0$ can be found in [3, 5]:
$\lambda=\sqrt{1+\frac{\lambda_{m}^{2}}{1+\varphi^{2}(1+\lambda_{m}^{2})}},$ (4)
where $\lambda_{m}$ is the ratio between synchrotron and betatron horizontal beam sizes:
$\lambda_{m}=\frac{D_{x}^{*}\sigma_{\delta}}{\sqrt{\varepsilon_{x}\beta_{x}^{*}}}.$ (5)
From Eq. (4) it follows that $\sigma_{W}$ should increase from 26.4 to 37.9 MeV for ZH4IP and from 20.2 to 41.8 MeV for ZH2IP, making the latter scheme less attractive. These values do not agree well with the simulation results presented in Table 3, which raises questions about this simulation.
A brief explanation is needed here. Simulation results can often differ from analytical formulas, since the former typically take into account nonlinear effects, beam-beam resonances, etc., which lead to perturbations of the beam distribution. However, the simulations described in [1] do not include any of these. There, the beams were represented ”as an ideal Gaussian distribution defined by the global optical parameters”. Therefore, the result should match the analytical formula, which was derived exactly for this distribution.
@@ 5 Vertical emittance
In Section 3.3, the authors reasonably noted that $\varepsilon_{y}$ was calculated with the assumed (initial) FCC-ee coupling ratio $\varepsilon_{y0}\big/\varepsilon_{x0}=0.2\%$. If $D_{x}^{*}\neq 0$, $\varepsilon_{x}$ will increase several times due to BS, and it would be too optimistic to assume that $\varepsilon_{y}$ will remain unchanged.
The excitation of $\varepsilon_{y}$ depends on two factors: 1) betatron coupling and 2) vertical dispersion (mainly in the arcs) which also arises due to coupling. An increase in $\varepsilon_{x}$ due to BS will lead to an increase in $\varepsilon_{y}$ due to the first factor, but not the second. To estimate this effect, it is necessary to know the ratio of the contributions of these two factors to $\varepsilon_{y0}$. However, this depends on the specific sources of X-Y coupling in the realistic magnetic lattice with errors, misalignments, etc., as well as on the methods used to correct them, and all these details are unknown for now. For a realistic assessment, it is best to use a set of simulations (with different random seeds) that take into account all the details mentioned above. If such a model is not available, then only approximate assumptions can be made.
For example, $\varepsilon_{x}$ increased about 3.7 times for ZH4IP. Based on the conservative assumption that the contribution of the first factor to the nominal $\varepsilon_{y0}$ is at least one-third, we will get a 1.9-fold increase in $\varepsilon_{y}$ due to BS. Accordingly, the luminosity will decrease by approximately 1.4 times, and this factor (with appropriate explanations) should be taken into account in the Table of parameters to make it more realistic.
@@ 6 Parameter optimization
The Abstract of [1] contains a phrase that is perceived as a promise that will be fulfilled: ”This measurement is significantly facilitated if the CM energy spread of $e^{+}e^{-}$ collisions can be reduced to a level comparable to the natural width of the Higgs boson, $\Gamma_{H}$ = 4.1 MeV, without substantial loss in luminosity”. And here is the result: $\sigma_{W}$ became $\sim$38 MeV, while the luminosity decreased by $\sim$2.1 times (taking into account the aforementioned factor of 1.4). Is this what we were aiming for?
As can be seen from Eq. (4), to get $\lambda\gg 1$ in a collision scheme with crossing angle and without crab crossing, we need to make $\varphi\ll 1$. To achieve this, $\sigma_{z}$ can be reduced (for example, $\sim$1.5 times) by increasing the RF voltage and $D_{x}^{*}$ should be significantly increased. The statement about a ”minimum bunch spacing of 25 ns” also raises questions. Perhaps this limitation exists at low energy (Z-pole, 45.6 GeV), where the total beam current is 3.5 times higher. However, at 62.5 GeV things become easier, so decreasing $N_{b}$ (along with increasing the number of bunches $n_{b}$) can be a useful option.
An adequate example of parameter optimization can be found in the CDR [5], where the emphasis was on achieving the required $\sigma_{W}$ and luminosity values. Of course, much has changed since then, including the perimeter, the number of IPs, the arc cell type at low energies and hence $\alpha_{C}$, $\varepsilon_{x}$, etc. Consequently, it will be necessary to re-optimize all the parameters, and then, which will be the most difficult part, implement this in the IR optics design.